OpenStudy (anonymous):

Convert the function into intercept form. Show your work. y=-x^2+5x+36

OpenStudy (anonymous):

\[ y=-x^2+5x+36 \rightarrow y=-(x^2-5x-36)\] \[\rightarrow y=-(x^2-9x+4x-36)\] \[\rightarrow y=-[x(x-9)+4(x-9)]\] \[\rightarrow y=-(x-9)(x+4)\] Hence y=-(x-9)(x+4) is the required intercept form of the given quadratic function. @adra_05

OpenStudy (anonymous):

@adra_05

OpenStudy (anonymous):

this was so much help thank you :)

OpenStudy (anonymous):

@adra_05 Medal please Click on best response button.